OpenStudy (anikate):

Fin the vertex and axis of symmetry of each quadratic. f(x)=-(x+3)(x+5)

OpenStudy (campbell_st):

the roots of your quadratic are at x = -3 and x = -5 the axis of symmetry is midway between the roots to find it is x = (-3 + -5)/2 x = - 4 is the line of symmetry In general for a quadratic \[ ax^2 + bx + c = 0\] then the axis of symmetry is \[x = \frac{-b}{2a}\] the vertex lies on the axis of symmetry. to find it you'll need to find f(-4) then the vertex will be at the ordered pair (-4, f(-4))